OpenStudy (anonymous):

Given the function g(x) = 4x − 5, find g(−3). −17 7 −7 2

OpenStudy (anonymous):

-17

OpenStudy (anonymous):

4(-3) - 5

OpenStudy (anonymous):

-12 - 5

OpenStudy (anonymous):

-17

OpenStudy (hoblos):

put (-3) in place of x g(−3)= 4(-3)-5 = -12-5=-17